OpenStudy (anonymous):

how to differentiate z=(1-2v+v^4)^-3/2

OpenStudy (anonymous):

you should use chain rule

OpenStudy (anonymous):

dz/dv=-3/2(1-2v+v^4)^(-3/2-1) * (-2+4v^3)

OpenStudy (anonymous):

oh? thanks for the hint :)

OpenStudy (anonymous):

like this let u = 1 - 2v + v^4 then \[\frac{ dz }{ du } u ^{\frac{ 3 }{ -2 }}*\frac{ du }{ dv }\]

OpenStudy (anonymous):

like this

OpenStudy (anonymous):

welcome..

OpenStudy (anonymous):

i come up with z'=3(1-2v^3)/square rt. of (1-2v+v4)^5 :)

OpenStudy (anonymous):

ur correct @doodlellatot

OpenStudy (anonymous):

thanks helping me :)

OpenStudy (anonymous):

go to calc101.com you will love it, it has working and everything